An empty swimming pool can be filled in 10 hours with the Honda engine, and the other engine will fill the same swimming pool in 19 hours. How long will it take to fill the pool if both engines are left open?
Answer by edjones(7569)

In 1 hr honda does 1/10 the job.
In 1 hr brand x does 1/19 the job.
1/10 + 1/19
19/190 + 10/190 = 29/190 of the job done in 1 hr working together.
190/29 = time in hrs it takes both of them to do the job.
.
Ed